OpenStudy (anonymous):

show that root 3+root 5 is irrational

OpenStudy (science0229):

Ok. First, if we can show that sqrt(3) and sqrt(5) are both irrational, then this can be easily proven, right?

OpenStudy (science0229):

I'll be using reductio ad absurdum

OpenStudy (science0229):

Assume that sqrt(3) is a rational number So, it can be written like this:\[\sqrt3=\frac{ p }{ q }\] where p and q are distinct natural numbers.

OpenStudy (science0229):

square both sides and get\[3=\frac{ p^2 }{ q^2 }\] \[p^2=3q^2\] So, p^2 is a multiple of 3.

OpenStudy (science0229):

That means that p has to be a multiple of 3 since p is a natural number. So, we can say that \[p=3k\] k is a natural number constant substitute that in the last equation to get \[(3k)^2=3q^2\] \[9k^2=3q^2\] \[q^2=3k^2\] So q^2 is also multiple of 3, which means that q is multiple of 3. So, p and q have a common factor of 3. But in the first place, we assumed that both of them are distinct, or without common factor. So the assumption that sqrt(3) is wrong.

OpenStudy (science0229):

Correction: Assumption that sqrt(3) is rational number is wrong. Similarly, we can prove that sqrt(5) is not rational number, either. Becuase both of them are irrational number, sqrt(3)+sqrt(5) is irrational number.